UARS Weekly Status Report

25 January 2002




********************************************************************

                UARS WEEKLY STATUS REPORT - 25 JAN 2002

********************************************************************

***********
SOLSTICE
***********

SOLSTICE continues to operate normally, taking solar data on all
available orbital tracking periods. We are prepared for closed loop 
SSPP tracking expected to begin on January 30th.

***********
SUSIM
***********

SUSIM completed daily mid resolution and MgII solar scans during
periods of predicted UARS telemetry.  We are awaiting the corrected
level 0 data from the PACOR-A system.  On a positive note, we have
been able to successfully download cataloged files from the former
CDHF.

***********
ACRIM
***********

Nothing received.

***********
PEM
***********

PEM continues to receive NMC data files from NOAA.  No problems have
been encountered.

An initial review of the PEM CICEL PROCs were conducted.  Comments have 
been submitted.  The converted STOL procedures are not acceptable for use at 
the current time.

PEM awaits support of the UARS Traceability Mission for the TIMED mission.

*Beginning of power orbit day (PEM activation): array current 5 amps, array 
     voltage 53.5 volts, and battery current 2.0 amps.  
 End of power orbit day (PEM deactivation): time in power orbit day 4 minutes,
     array current <9 amps, and battery currents <0.2 amps.
 Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.

***********
HALOE 
***********

HALOE continues to acquire its data as scheduled.  The processing of scans
acquired since November will commence shortly on the Remote Analysis
Computer (RAC) at Langley.  Those data will be posted initially through
the HALOE Website.

One problem remains with our present adaptation of the operational
software and the processing from the Level 0 files. Whenever a HALOE 
sunrise or sunset scan event crosses the day boundary, the UARS Orbit 
and Attitude Services (UOAS) software must open and access two definitive 
orbit files because these are 24hr files.  It accesses the current processing 
day file but bombs when accessing the next day file.

It is worth noting that we have never before used the UOAS software
outside of the CDHF computer environment.  At CDHF the processing is done 
either in an 'Engineering Mode' or a 'Production Mode', and in both cases 
the HALOE software is linked with a 'blackbox' of libraries which consists of 
the UOAS and UCSS, among other things.  What is necessary to process the data
at Langley is to adapt these libraries to perform the same functions on
our RAC.  We have always had the UCSS library on our RAC (though for
terminology confusion it is called an RSS library locally) or we would not 
have been able to run Level 1 or Level 2 at Langley.  To do
the Level 0 processing at Langley, we now have to add the UOAS library
which has dependence on our RAC RSS library; the problem is most likely
with this new dependence.

If anyone from the other instrument teams can offer some insight about how to 
overcome this problem, please contact Lance Deaver (or Ellis Remsberg) of the 
HALOE Project.  In the meantime we will perform a 'work around', losing at 
most one event on those days when it straddles the day boundary.  
This represents a loss of one scan out of, perhaps, 26 for that day.  
We anticipate being able to process data in support of the Traceability Option 
for the TIMED and SAGE III satellite experiments in the coming weeks.

***********
MLS
***********

Nothing received.

***********
HRDI
***********

Nothing received.

***********
WINDII
***********

Nothing received.

***********
MPG
***********

The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop 
capabilities for the Science Traceability Mission.

Daily processing of instrument Daily Activity Plans 
continues with no problems.

***********
CDHF
***********

Nothing received.
                  
***********
FOT
***********

                            REPORTING PERIOD

This  report covers 12 January 2002 (Orbit 56523, GMT Day 012) through
18 January 2002 (Orbit 56627, GMT Day 018).

                          SPACECRAFT OPERATION

The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and   is  performing  nominally  under  two-battery  operations.   The
instrument operational changes were:


                     INSTRUMENT OPERATIONAL CHANGES


        Instrument    Time          Orbit     Comment
        WINDII        014/2325      56567     Power OFF


The  beta  angle ranged between the angles of 73.67 degrees  to  68.63
degrees  for  this  report period.  The beta angle is  now  decreasing
towards a minimum angle of 0 degrees on 03 February 2002 (DOY 034).

Spacecraft  battery  1,  2,  and 3 performance  monitoring  continues.
Battery 1 remains OFF the charge relay.  Battery 2 and Battery  3  are
currently maintaining greater than 26.72 V end-of-night (EON) load bus
voltage.   Battery  temperatures are stable with a  temperature  delta
between Battery 2 and Battery 3 of 1.25 to 0.94 degrees C.  Additional
battery performance data for each battery on 22 December 2001  and  28
December 2001 are provided in Appendix D.

The  clock  error  ranged from 11.6 to 13.0 msec during  this  report
period.  Clock rate adjustments are listed in Appendix C.

All  instrument  and subsystem engineers were advised on  the  initial
excursion  of ALERTS or Out-Of-Limits conditions detected by  the  FOT
for  this report.  A summary is listed below and the details of  ALERT
and Out-Of-Limits occurrences are listed in Appendix B.

     OUT-OF-LIMITS  =    00
     ALERTS         =    00

The  following real-time command sheets were executed for  the  listed
operational element(s) during this report period (see Appendix A).

    FOT            =    8
    PEM            =    1
    WINDII         =    1

UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.

     FOT            =    2

One (1) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).

TRANSPONDER FREQUENCIES:
    Transponder A:  01/333  - 2287.496855
    Transponder B:  01/324  - 2287.498688


                         GROUND SYSTEM OPERATION


Two  (2)  operational support problems occurred this period.   Anomaly
Reports are included in Appendix E.

The  data  loss  calculations are no longer  being  supplied  by  Data
Capture Facility (DCF). This service should resume soon.



                        OTHER SIGNIFICANT EVENTS


                        UARS SATELLITE OPERATIONS


More  realtime  tests were executed on the Eclipse system  this  week.
The  FOT now routinely takes two passes a day with the Eclipse system.
These  passes  are merely telemetry passes with only  NOOPs  sent  for
commands.  All loading and other spacecraft commanding is still  being
done  with  the  legacy  AP  system.  A new delivery  of  the  Eclipse
software  is  expected on 25 January 2002, which will include  command
load  capability.   This will be the focus of  testing  for  the  week
subsequent to that date (with the simulator).




                                MEETINGS

The  FOT  attended  a  UARS Ground System Rehost  Meeting  on  Monday,
January 14, 2002.
The  FOT  attended a UARS Eclipse Working Meeting on Thursday, January
17, 2002.
The FOT attended a UARS Flight Software Meeting on Friday, January 18,
2002.

                        FUTURE SIGNIFICANT EVENTS


The  next  UARS Yaw slew is scheduled for 03 February 2002 (DOY  034).
This  maneuver will be a Reverse to Forward Flight maneuver  (Southern
Hemisphere viewing).


 Attachments:    Appendix A, Real-time Command Sheets
                 Appendix B, Out-Of-Limits & ALERT Occurrences
                 Appendix C, Clock & Solar Array Rate Corrections
                 Appendix D, UARS Battery Performance Data
                 Appendix E, UARS Anomaly Reports
                               APPENDIX A
                        REAL-TIME COMMAND SHEETS
                       Week ending 18 January 2002

 ORBIT   AOS        NO.      SUBSYS      REASON
 -----   ---        ---      ------      ------
 56547   013/1603   02-010   OBC         Clk Adj fr 41.2 to 20.6 msec/day
 56563   014/1644   02-011   CDH         Eclipse cmd test
 56567   014/2325   02-012   WINDII      Power OFF
 56574   015/1046   02-013   OBC         Clk Adj fr 20.6 to 41.2 msec/day
 56578   015/1644   02-015   CDH         Eclipse cmd test
 56579   015/1826   02-014   PEM         ZEPS/HEPS Fulltime
 56580   015/2005   02-015   CDH         Eclipse Cmd Test
 56617   018/0717   02-016   OBC/CLK     Clk Adj fr 41.2 to 20.6 msec/day
 56623   018/1643   02-018   CDH         ECLIPSE TEST CMDS
 56625   018/2002   02-018   CDH         ECLIPSE TEST CMDS

                               APPENDIX B
                   OUT-OF-LIMITS & ALERTS OCCURRENCES
                       Week ending 18 January 2002

 The following Out-Of-Limits occurrences were detected:

 MNEMONIC       STATE   DESCRIPTION                ORBITS
 --------       -----   -----------                ------
 -None this report period.


 The following ALERT occurrences were detected:

 MNEMONIC     DESCRIPTION                          ORBITS
 --------     -----------                          ------
 -None this report period.

 ______________________________________________________________________

                               APPENDIX C
                  CLOCK & SOLAR ARRAY RATE CORRECTIONS
                       Week ending 18 January 2002

                      CLOCK DRIFT RATE ADJUSTMENTS

 ORBIT          TIME                   CLOCK RATE CHANGE
 -----          -----                  -----------------
 56547          013/1603               Clk Adj fr 41.2 to 20.6 msec/day
 56574          015/1046               Clk Adj fr 20.6 to 41.2 msec/day
 56617          018/0717               Clk Adj fr 41.2 to 20.6 msec/day


                               APPENDIX D
                      UARS BATTERY PERFORMANCE DATA
                       Week ending 18 January 2002


   12 January 2002 - GMT Day 012 Beta = 73.67 deg, SA Pos = 269 deg

      DIFF V (mV)  TEMP (C)     CURR (amp)  EON V    AVE C/D  AVE
 BAT  MAX/MIN      MAX/MIN      MAX/MIN     MIN      RATIO    DOD(%)
 ---  -------      --------     ----------  -----    -------  ------
 1  -504.0/-672.0  -2.85/-3.77  +0.0/ -0.0  10.72    0.000      0.00
 2   000.0/-224.0  +2.42/+2.20  +16.0/-12.8 27.52    1.050     13.17
 3  +224.0/-112.0  +1.88/+1.23  +15.6/-12.8 27.52    1.040     12.93


   18 January 2002 - GMT Day 018 Beta = 68.63 deg, SA Pos = 269 deg

      DIFF V (mV)   TEMP (C)     CURR (amp)   EON V    AVE C/D  AVE
 BAT  MAX/MIN       MAX/MIN      MAX/MIN      MIN      RATIO    DOD(%)
 ---  -------      ----------   ----------    -----    -------  ------
 1  -560.0/-728.0   -3.16/-3.77  +0.0/ -0.0   10.72    0.000    0.00
 2  +056.0/-224.0   +3.17/+2.52  +18.8/-14.4  26.88    1.040   15.85
 3  +224.0/-112.0   +2.52/+1.88  +18.4/-14.4  26.72    1.040   15.38

       V/T LEVEL CHANGES
 ORB#          TIME       LVL FR-TO
 -----         ----       ---------
 -no changes    V/T 5




                               APPENDIX E
                          UARS ANOMALY REPORTS
                       Week ending 18 January 2002

                      NEW ANOMALY REPORTS GENERATED
 AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
 ------    ------    -----     ----      ---       -----------------
 02-001    FOT       56577     015/1600  N/A       TAC/AP Crashed
 02-002    FOT       56592     016/1559  14170     Generic Late
 Acquisition

                 ANOMALY CLOSURE INFORMATION RECEIVED
 AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
 ------    ------    -----     ----      ---       -----------------
 02-001    FOT       56577     015/1600  N/A       TAC/AP Crashed

                   ANOMALY INVESTIGATION REPORTS GENERATED
 AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
 ------    ------    -----     ----      ---       -----------------
 -None this report period.

                          OPEN ANOMALY REPORTS
 AR NO.    SUBSYS    ORBIT     TIME      TTR       PROBLEM/DATA LOSS
 ------    ------    -----     ----      ---       -----------------
 02-002    FOT       see list  Jan 02    14170     Generic Late Acquisition